3(x + 6)= 4x I need to solve this and show a check.
Answer by richard1234(7193)

Multiplying the left side we get:
3x + 18 = 4x --> x = 18
It's easy to check by plugging x = 18 into the equation:
3(18 + 6) = 4(18)
72 = 72